An investor buys a total of 360 tickets for CONCERT. The price of adult ticket is $35, while the price for children's ticket is $45. The investor spends a total of $15000. How many tickets for adult and for children did he buy?
Answer by VFBundy(438)

# of Adult tickets = A
# of Children's tickets = C

Set up two equations:

A + C = 360
35A + 45C = 15000

Multiply the first equation by -45:

-45(A + C) = -45(360)

-45A - 45C = -16200

Here are your two equations again:

-45A - 45C = -16200
35A + 45C = 15000

Add the equations together:

-10A = -1200

Solve for A:

A = 120

Go back to the original equations we set up. The first one is:

A + C = 360

Plug in A = 120:

120 + C = 360

Solve for C:

C = 240

# of Adult tickets = A = 120
# of Children's tickets = C = 240